In Problems 7–10, construct a confidence interval for p - P2 (it the given level of confidence. 7. x1 = 368, n, = 541. X2 = 4\
Answer #1

p1= x1/n1 = 368/541= 0.6802

p2= x2/n2 = 421/593 = 0.7099

So, 90% CI for the difference in the population proportion is given as:

p1- p2 +- z 0.05*√p1*(1-p1)/n1 + p2*(1-p2)/n2

= -0.1017 +- 1.645*0.0281

= (-0.148, -0.056)

We are 90% confident that the true difference between p1 and p2 lies in the above interval.